OpenStudy (jennychan12):

Suppose f is one-to-one, f(7) = 8, and f'(7) = 8. Find a) f^-1(3) and b) f^-1'(3). I'm kinda stuck on how to find part b). Also check my work for part a) ?

OpenStudy (anonymous):

does this problem come with a graph?

OpenStudy (jennychan12):

y-8 = 8(x-7) y = 8x+48 so f^-1(x) = x = 8y - 48 so x = 8y - 48 so 8y = x+48 so y = (x+48)/8 so f^-1(3) = 51/8 = 6.375

OpenStudy (jennychan12):

no.

OpenStudy (anonymous):

ok you found \[y=\frac{ x+48 }{ 8 }\] for the inverse function right? do you know how to take the derivatives yet?

OpenStudy (jennychan12):

yeah. -_- OMG. it's that easy... my teacher showed us a TOTALLY confusing way to do it. so i guess i just got too confused. so i think i just answered my own question. how sad. :(

OpenStudy (anonymous):

haha it's ok :) glad you figured out how to do it!

OpenStudy (jennychan12):

is there a way to do this without finding the inverse function?

OpenStudy (anonymous):

yes

OpenStudy (jennychan12):

would you just take the derivative of the original function and flip it? and that would be the derivative of the inverse? idk just a guess.

OpenStudy (anonymous):

but i have a question how do you know the function is a line?

OpenStudy (anonymous):

it doesn't say anything about it being a line if it was a line, they would say \(f'(x)=8\) not \(f'(7)=8\)

OpenStudy (jennychan12):

lol i don't. i just assumed. ohhhhh now i see it.

OpenStudy (anonymous):

in fact there must be some other part missing in the question knowing that \(f(8)=7\) and that \(f'(8)=7\) tells you nothing about this function at 3 or its inverse at 3

OpenStudy (jennychan12):

it just says find inverse and inverse derivative at 3. that's all that's given; what i typed.

OpenStudy (anonymous):

can you take a picture of the problem?

OpenStudy (jennychan12):

lol my teacher got it from a random website and she posted the problem on her website. and i just copy and pasted whatever she had. THERE IS NO GRAPH. Suppose f is one-to-one, \[f(7) = 8\] \[f'(7) = 8\] Find a) \[f^{-1}(3) \] b) \[f'^{-1}(3)\] I just wrote it in a better format. i took a picture anyways. it's uploading.

OpenStudy (anonymous):

what if she was just trolling you when the real answer would be "cannot conclude answer without enough information"

OpenStudy (jennychan12):

ignore my work and other stuff stuff. it's #2.

OpenStudy (anonymous):

i feel like she copied the problem wrong. i don't know any way to acquire this without assuming the equation was linear :( sorry